OpenStudy (anonymous):

determination the number k so that the polynomial x^3+kx^2-kx+9 gets a factor x+3

OpenStudy (lgbasallote):

are you familiar with remainder theorem?

OpenStudy (lgbasallote):

psst @kalemale ..you there?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

does not say anything about that in my book

OpenStudy (lgbasallote):

well that's a bummer...how about long division? familar?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

let's go with that...

OpenStudy (lgbasallote):

x^2 + (k-3)x + (-4k + 9) ______________________ x + 3 | x^3 + kx^2 -kx + 9 x^3 + 3x^2 =============== (k-3)x^2 - kx (k-3)x^2+(3k-9)x ================ (-4k+9)x + 9 (-4k+9)x +(-12k +27) since it's a factor, the remainder should be 0 so 9 - (-12k+27) = 0 9 + 12k -27 = 0 -18 + 12k = 0 12k = 18 k = 18/12 k = 3/2 do you get that?

OpenStudy (anonymous):

ye I see now can you do one with the remainder therom? It seems much easier. Thanks for all the help!

OpenStudy (lgbasallote):

now i'll show you the remainder theorem to prove im right since x + 3 is a factor, x = -3 is a root therefore p(-3) = 0 P(-3) = (-3)^3 + k(-3)^2 - k(-3) + 9 -27 + 9k + 3k + 9 = 0 12k - 18 = 0 12k = 18 k = 18/12 k = 3/2

OpenStudy (anonymous):

Thanks!!!! =)

OpenStudy (lgbasallote):

welcome ^_^